### v2

# Nef Cones of the Hilbert Schemes of Points on Generalized Cayley K3 Surfaces

Abstract.
We study the nef cones and fundamental domains of Hilbert schemes of points on the Cayley K3 surface $S$ and its generalizations $S_{a}$. For the Hilbert square $S^{[2]}$, we explicitly compute the nef cone and describe a fundamental domain using the automorphisms of $S^{[2]}$ and lattice-theoretic methods. For higher Hilbert schemes $S_{a}^{[n]}$, we determine the nef cones using Bridgeland stability methods that identify the contracted curves defining walls and the divisors generating the extremal rays.
Key words and phrases:
Hilbert scheme, K3 surface, nef cone
2020 Mathematics Subject Classification:
14J28, 14C05
1. Introduction
Hilbert schemes of points on K3 surfaces are classical and important examples of higher-dimensional holomorphic symplectic manifolds. By Fogarty [12], the Hilbert scheme $S^{[n]}$ of length-$n$ subschemes of a smooth projective surface $S$ over $\mathbb{C}$ is smooth, irreducible, and projective, and has dimension $2n$. Fogarty [13] later computed its Picard group, proving that
$\operatorname{Pic}(S^{[n]})\cong\operatorname{Pic}(S)\oplus\mathbb{Z}e,$
where $e=\tfrac{1}{2}B$ denotes half of the class of the exceptional divisor $B$ of the Hilbert–Chow morphism. When $S$ is a K3 surface, Beauville [7] showed that $S^{[n]}$ admits a holomorphic symplectic structure, thereby providing higher-dimensional examples of compact hyperkähler manifolds. In the same work, he introduced a quadratic form on $H^{2}(S^{[n]},\mathbb{Z})$. Fujiki [14] later established its cohomological properties, leading to what is now called the Beauville–Bogomolov–Fujiki (BBF) form, which underlies the intersection theory and birational geometry of these varieties.
One of the main problems in the study of hyperkähler varieties is to describe the nef cone. For K3 surfaces, Sterk [27] proved that the nef cone is determined by walls orthogonal to (2)(-2)-curves, and that the automorphism group acts on it with a rational polyhedral fundamental domain. This was the first verification of what later became the Morrison–Kawamata cone conjecture [23, 18], which predicts the existence of such a fundamental domain for Calabi–Yau and hyperkähler manifolds in general.
Markman [21, 22] described the nef and movable cones from a monodromy-theoretic perspective. Amerik and Verbitsky [1] introduced the notion of monodromy birationally minimal (MBM) classes and proved the Morrison–Kawamata cone conjecture for all hyperkähler manifolds.
The nef cones of Hilbert schemes of points on K3 surfaces have also been studied through explicit lattice-theoretic and numerical approaches. Hassett and Tschinkel [15, 16] described the structure of the ample and movable cones. Bayer, Hassett, and Tschinkel [3] extended this analysis to projective K3[n]-type varieties.
Bridgeland stability conditions provide a new conceptual approach to studying these problems. Bayer and Macrì first used wall-crossing to describe the nef and movable cones of moduli spaces of sheaves on K3 surfaces [5], and subsequently established the Positivity Lemma [4], which produces nef divisors from stability conditions and identifies the curves orthogonal to them. Liu [20] gave a general formulation of the local Bayer–Macrì decomposition of nef divisors. The wall–chamber structure for Hilbert schemes of points was analyzed in detail in [8].
The Cayley K3 surface is a concrete example of a K3 surface that is not a Mori dream surface. It is a determinantal quartic surface $S\subset\mathbb{P}^{3}$ whose Picard lattice has rank two and intersection form
$\begin{pmatrix}4&2\\ 2&-4\end{pmatrix}$.
This surface was already considered by Cayley in the nineteenth century [11] and later studied by Oguiso [25]. More recently, Lee [19] studied the automorphism group of the Hilbert square $S^{[2]}$ of the Cayley K3 surface.
In this paper, we study the structure of the nef cones and fundamental domains for Hilbert schemes of points on a family of K3 surfaces $S_{a}$. We fix a basis $\{h_{1},h_{2}\}$ of $N^{1}(S_{a})$ such that the intersection form on $N^{1}(S_{a})$ is represented by
$\begin{pmatrix}4&2a\\ 2a&-4\end{pmatrix}$,
where $a\geq 1$ is an integer. The family of K3 surfaces $S_{a}$ generalizes the Cayley K3 surface, which corresponds to the case $a=1$.
Section 3 presents a computation of the nef and effective cones for the special case of the Hilbert square $S^{[2]}$ of the Cayley K3 surface, using lattice-theoretic calculations involving the BBF form. In particular, we prove the following:
Theorem 1.1.
(=Theorem 3.2) Every wall of the nef cone of the Hilbert square of the Cayley K3 surface $S_{1}$ is given by $(\iota\cdot e)^{\perp}$, where $\iota\in\operatorname{Aut}(S^{[2]})$.
In Section 4 we construct a rational polyhedral fundamental domain for the action of $\operatorname{Aut}(S_{1}^{[2]})\cong\mathbb{Z}_{2}*\mathbb{Z}_{2}*\mathbb{Z}_{2}$ in Theorem 4.1. Finally, Section 5 studies the more general family of K3 surfaces $S_{a}$ and higher Hilbert schemes $S_{a}^{[n]}$ for sufficiently large $n$. Theorem 5.1 and Corollary 5.2 provide a complete description of their nef and Mori cones (i.e., the closure of the cone of curves), using the Positivity Lemma to identify the corresponding extremal nef divisors and contracted curves.
Acknowledgments
The author would like to thank Yongnam Lee for detailed comments on the draft of this paper, and Wanmin Liu for helpful comments on the references. This work was supported by the Institute for Basic Science (IBS-R032-D1).
2. Preliminaries
2.1. K3 surfaces and their Hilbert schemes
We work over $\mathbb{C}$. A K3 surface is a smooth, projective, simply connected surface $S$ with a trivial canonical bundle. For a reference on the theory of K3 surfaces and their lattices, see [17].
The second cohomology lattice of a K3 surface is isometric to the K3 lattice
$\Lambda_{\mathrm{K3}}\;\cong\;U^{\oplus 3}\oplus E_{8}(-1)^{\oplus 2},$
where $U$ denotes the hyperbolic plane (the even unimodular lattice of signature $(1,1)$) and $E_{8}(-1)$ the negative definite $E_{8}$ lattice.
By Nikulin’s theorem [24], every even hyperbolic lattice $L$ of rank at most $20$ that satisfies the usual discriminant-form condition (i.e. the discriminant form of $L$ embeds into that of $\Lambda_{\mathrm{K3}}$) admits a primitive embedding into the K3 lattice. In particular, for each integer $a\geq 1$ there exists a projective K3 surface whose Picard lattice is
$\begin{pmatrix}4&2a\\ 2a&-4\end{pmatrix}$,
with respect to a basis $(h_{1},h_{2})$. We denote this surface by $S_{a}$.
The Hilbert scheme of $n$ points is the moduli space parameterizing zero-dimensional closed subschemes of length $n$. The Hilbert scheme $S^{[n]}$ of $n$ points on a K3 surface is an irreducible holomorphic symplectic manifold [7]. Since $\operatorname{Pic}(S)$ embeds into $\operatorname{Pic}\!\left(S^{[n]}\right)$ via $D\mapsto D^{[n]}$, we identify $h_{i}$ with $(h_{i})^{[n]}$.
There is a natural quadratic form on $H^{2}(X,\mathbb{Z})$ of irreducible holomorphic symplectic manifolds:
Definition 2.1 (Beauville–Bogomolov–Fujiki form [7, 14]).
Let $X$ be an irreducible holomorphic symplectic manifold of complex dimension $2n$. There exists a primitive integral quadratic form
$q_{X}\colon H^{2}(X,\mathbb{Z})\longrightarrow\mathbb{Z}$
of signature $(3,b_{2}(X)-3)$, which is called the Beauville–Bogomolov–Fujiki (BBF) form. The associated bilinear form is defined by
$(D,D^{\prime})=\tfrac{1}{2}\big(q_{X}(D+D^{\prime})-q_{X}(D)-q_{X}(D^{\prime})\big),\qquad D,D^{\prime}\in H^{2}(X,\mathbb{Z}).$
It induces a natural identification $H^{2}(X,\mathbb{Z})\cong H_{2}(X,\mathbb{Z})$, so that divisor classes and curve classes can be treated in the same lattice.
Let $S$ be a K3 surface and $X=S^{[n]}$ its Hilbert scheme of length-$n$ subschemes. Then there is an orthogonal decomposition with respect to the BBF form
$H^{2}(S^{[n]},\mathbb{Z})\;\cong\;H^{2}(S,\mathbb{Z})\;\oplus\;\mathbb{Z}\,e,$
where $e=\tfrac{1}{2}B$ denotes the class of half the exceptional divisor $B$ of the Hilbert–Chow morphism $S^{[n]}\to S^{(n)}$. Under the identification, $q_{X}$ restricts to the usual cup product on $H^{2}(S,\mathbb{Z})$ and takes the value $q_{X}(e)=-2(n-1)$.
We denote by $C_{0}$ the class of a curve contracted by the Hilbert–Chow morphism. If $C\subset S$ is a curve that admits a $g^{1}_{n}$, then the fibers of this map induce a rational curve $C_{[n]}$ on $S^{[n]}$. The intersection numbers are given as follows:
$\begin{array}[]{c|cc}&D^{[n]}&B\\ \hline\cr C_{[n]}&C\cdot D&2g(C)-2+2n\\ C_{0}&0&-2\end{array}$
Note that for the Hilbert scheme $S_{a}^{[n]}$, the BBF form on $\operatorname{NS}(S_{a}^{[n]})$ is given in the basis $(h_{1},e,h_{2})$ by
$q(xh_{1}+ye+zh_{2})=4x^{2}+4axz-4z^{2}-2(n-1)y^{2}.$
2.2. Cones of hyperkähler manifolds
Let $X$ be an irreducible holomorphic symplectic (IHS) manifold, that is, a simply connected compact Kähler manifold with $H^{2,0}(X)=\mathbb{C}\sigma$ spanned by a holomorphic symplectic form $\sigma$. If $X$ is deformation equivalent to the Hilbert scheme $S^{[n]}$ of a K3 surface $S$, we say that $X$ is of K3[n]-type.
Inside $H^{1,1}(X,\mathbb{R})$ one defines the positive cone
$\operatorname{Pos}(X)=\{x\in H^{1,1}(X,\mathbb{R})\mid q_{X}(x)>0\},$
which contains the ample cone $\operatorname{Amp}(X)$.
The structure of these cones for manifolds of $\mathrm{K3}^{[n]}$-type is controlled by certain distinguished divisor classes, called wall divisors. The following numerical criterion, due to Hassett and Tschinkel, characterizes them in the case $n=2$. It will serve as the starting point for our later computations of nef cones.
Theorem 2.2 ([16]).
Let $X$ be a projective algebraic variety deformation equivalent to the Hilbert square of a K3 surface. Fix an ample divisor $g$ on $X$. A divisor $h$ on $X$ is ample if $(h,\rho)>0$ for every divisor class $\rho$ satisfying $(g,\rho)>0$ and either
  1. (1)
    $(\rho,\rho)\geq-2$ , or
  2. (2)
    $(\rho,\rho)=-10$ and $(\rho,H^{2}(X,\mathbb{Z}))=2\mathbb{Z}$.
2.3. Cone conjecture
We recall the Morrison–Kawamata cone conjecture in the setting of projective hyperkähler manifolds.
Conjecture 2.3 (Morrison–Kawamata).
Let $X$ be a projective hyperkähler manifold. Then the action of $\operatorname{Aut}(X)$ on the nef-effective cone $\operatorname{Nef}^{+}(X)$ admits a rational polyhedral fundamental domain.
When the cone conjecture holds, the nef-effective cone
${\rm Nef}^{e}(S^{[2]}):={\rm Nef}(S^{[2]})\cap{\rm Eff}(S^{[2]})$
coincides with the rational hull of the nef cone
$\operatorname{Nef}^{+}(S^{[2]})={\rm conv}\big(\operatorname{Nef}(S^{[2]})\cap H^{1,1}_{\mathbb{Q}}(S^{[2]})\big)\subset H^{1,1}_{\mathbb{Q}}(S^{[2]})\otimes_{\mathbb{Q}}\mathbb{R}.$
Since every wall of the nef cone is rational, and since $q(x,y,z)=0$ admits no nonzero rational solution, we have
$\operatorname{Nef}^{+}(S^{[2]})=\operatorname{Nef}^{e}(S^{[2]})=\operatorname{Nef}(S^{[2]})\cap\operatorname{Pos}(S^{[2]}).$
A version of the conjecture was proved by Amerik and Verbitsky [2] for simple hyperkähler manifolds with $b_{2}(X)\neq 5$. Their result establishes the conjecture in an analytic setting and, in the projective case, implies the classical form; in particular, the existence of a rational polyhedral fundamental domain for the action of the automorphism group on the nef cone.
For the Hilbert square of the Cayley K3 surface, the automorphism group acts discretely on $\operatorname{NS}(S^{[2]})_{\mathbb{R}}$, preserving the BBF form. By [19], it is generated by three automorphisms and is isomorphic to the free product $\operatorname{Aut}(S^{[2]})\cong\mathbb{Z}_{2}*\mathbb{Z}_{2}*\mathbb{Z}_{2}$. This action admits a rational polyhedral fundamental domain, whose explicit lattice-theoretic construction will be given in Section 4.
2.4. Beauville involution
Beauville involutions arise as natural automorphisms of Hilbert squares of K3 surfaces. Beauville [6] showed that quartic K3 surfaces without lines admit involutions of this type. For the Cayley K3 surface, these involutions are essential in describing the automorphism group and the geometry of its nef cone.
Definition 2.4 (Beauville involution).
Let $S$ be a K3 surface and $X=S^{[2]}$. Suppose $D\in H^{2}(X,\mathbb{Z})$ is a primitive class with $q(D)=2$. The Beauville involution is the lattice isometry
$\iota(x)=-x+\frac{2(x,D)}{q(D)}D,$
which fixes $D$ and acts as 1-1 on its orthogonal complement $D^{\perp}\subset H^{2}(X,\mathbb{Z})$.
Remark 2.5.
If $D$ is of the form HeH-e with $H$ very ample of square $4$ on $S$, then $S$ embeds as a quartic surface in $\mathbb{P}^{3}$. In this case the lattice involution $\iota$ admits a geometric description: for two general points $p,q\in S$, the line $\overline{pq}\subset\mathbb{P}^{3}$ meets $S$ in two further points $p^{\prime},q^{\prime}$, and the correspondence $(p,q)\mapsto(p^{\prime},q^{\prime})$ induces the same involution on $S^{[2]}$. Moreover, if the quartic $S$ contains no line, this correspondence is regular everywhere, and $\iota$ is an automorphism rather than a birational involution [6].
2.5. Bridgeland stability conditions
We briefly recall the basic notions of stability conditions. Bridgeland [9] introduced the general notion of a stability condition on a triangulated category. A stability condition is a pair $\sigma=(Z,\mathcal{P})$ consisting of a group homomorphism
$Z\colon K(\mathcal{D})\longrightarrow\mathbb{C},$
called the central charge, together with a slicing $\mathcal{P}$ of the category $\mathcal{D}$, such that $\sigma$ satisfies the support property. The space $\mathrm{Stab}(\mathcal{D})$ of all stability conditions carries the structure of a complex manifold, with local coordinates given by the central charge.
In the case of K3 surfaces, Bridgeland [10] proved the existence of stability conditions and described the distinguished connected component of geometric stability conditions. These can be parametrized by pairs $(\beta,\omega)\in\operatorname{NS}(S)_{\mathbb{R}}\times\operatorname{Amp}(S)$, where $\beta$ is a real divisor class and $\omega$ is an ample class. The associated central charge on a Mukai vector $v\in K(S)$ is given by
$Z_{\beta,\omega}(v)=-\langle e^{i\omega+\beta},v\rangle,$
where $\langle\;,\;\rangle$ denotes the Mukai pairing.
Given a K3 surface $S$, a primitive Mukai vector $v$, and a stability condition $\sigma$, we denote by $M_{\sigma}(v)$ the moduli space of $\sigma$-stable objects with Mukai vector $v$. Whenever this space is nonempty and smooth, it is an irreducible holomorphic symplectic variety.
To every stability condition $\sigma$ one can further associate a numerical divisor class $D_{\sigma}\in\operatorname{NS}(M_{\sigma}(v))$, constructed using the universal family and the Fourier–Mukai transform (see [5, Prop-Def. 3.2] for details).
Theorem 2.6 (Positivity Lemma [5, Thm. 1.1, Thm. 1.4]).
Let $S$ be a K3 surface, $v$ a primitive Mukai vector, and $\sigma$ a stability condition. Then the associated divisor class $D_{\sigma}$ on $M_{\sigma}(v)$ is nef. If $\sigma$ lies on a wall, then $D_{\sigma}$ is semiample and induces a contraction
$\phi_{\sigma}:M_{\sigma}(v)\to Y,$
and one has $D_{\sigma}\cdot C=0$ if and only if $C$ is contracted by $\phi_{\sigma}$.
These divisors are closely connected to the birational geometry of the moduli space. If a curve has intersection number zero with $D_{\sigma}$, then it is contracted by the morphism associated to $D_{\sigma}$. Such contractions arise when the stability condition lies on a wall, linking wall-crossing phenomena with the structure of the nef and Mori cones of $M_{\sigma}(v)$.
We begin by recalling that a general description of Bayer–Macrì line bundles on Bridgeland moduli spaces over surfaces is provided by the local Bayer–Macrì decomposition established by Liu [20]. In the present paper, we focus on the case of the Hilbert scheme $X=S^{[n]}$, corresponding to the Mukai vector v=(1,0,1n)v=(1,0,1-n); for later use, we adopt the description specialized to this setting, which appeared later in [8].
Theorem 2.7 (Walls in the $(H,D)$-slice [8, Lem. 3.3]).
In the $(H,D)$-slice, consider the numerical wall $W$ for $S^{[n]}$ defined by a rank-$1$ torsion-free sheaf
$F=I_{Z^{\prime}}(-L),$
where $Z^{\prime}$ is a zero-dimensional subscheme of length $w$ and $L$ is an effective divisor. The center $(s_{W},0)$ of this wall is given by
$s_{W}=-\,\frac{2(n-w)+L^{2}+2(D\cdot L)}{2(H\cdot L)}.$
Proposition 2.8 (Gieseker wall [8, Prop. 3.5]).
Assume $2n>D^{2}$, and suppose that the subobject giving the Gieseker wall for $S^{[n]}$ in the $(H,D)$-slice is a line bundle. Then the Gieseker wall is computed by $\mathcal{O}_{S}(-L)$, where
$L\in\operatorname{CrDiv}(H,D):=\bigl\{\,F\in\operatorname{Pic}(S)\mid F\text{ is effective and }H\cdot F<H\cdot(-D)\,\bigr\}\cup\{-D\}$
is chosen so that the numerical wall defined by $\mathcal{O}_{S}(-L)$ is as large as possible.
Proposition 2.9 (cf. [20, Lemma 4.8]; see also [8, Prop. 3.8]).
If $S$ has irregularity 0 and v=(1,0,1n)v=(1,0,1-n) is the Mukai vector for $S^{[n]}$, then the divisor $D_{\sigma}$ associated to the stability condition $\sigma$ is a multiple of
$\frac{1}{2}K_{S^{[n]}}-s_{W}H^{[n]}-D^{[n]}-e.$
These results show that the wall structure for $S^{[n]}$ is controlled by effective divisors on $S$, and that the nef and Mori cones of $S^{[n]}$ can be described by explicit lattice-theoretic conditions in $H^{2}(S^{[n]},\mathbb{Z})$, which are used throughout the arguments in Section 5.
2.6. Generalized Fibonacci sequences
For later lattice calculations, we recall the generalized Fibonacci recursion. Fix an integer $a\geq 1$ and define the $a$-generalized Fibonacci sequence $(a_{m})_{m\in\mathbb{Z}}$ by
$a_{0}=0,\quad a_{1}=1,\quad a_{m+2}=a\,a_{m+1}+a_{m}.$
For $a=1$ this is the classical Fibonacci sequence.
Let $\alpha$ and $\beta$ denote the roots of the equation $r^{2}-ar-1=0$:
$\alpha=\frac{a-\sqrt{a^{2}+4}}{2},\qquad\beta=\frac{a+\sqrt{a^{2}+4}}{2}.$
They satisfy $\alpha+\beta=a$ and $\alpha\beta=-1$.
Proposition 2.10.
The sequence $(a_{m})$ satisfies the following properties.
  1. (1)
    (Negative indices) For $m\geq 1$,
    $a_{-m}=(-1)^{m+1}a_{m}.$
  2. (2)
    (Addition and subtraction formulas) For all $m,n\in\mathbb{Z}$,
    $a_{m+n}=a_{m}a_{n+1}+a_{m-1}a_{n},\qquad a_{m-n}=a_{m}a_{n-1}-a_{m-1}a_{n}.$
  3. (3)
    (d’Ocagne identity) For all $m,n\in\mathbb{Z}$,
    $a_{m}a_{n+1}-a_{m+1}a_{n}=(-1)^{n}a_{m-n}.$
3. The nef cone of $S_{1}^{[2]}$
Let $a=1$ and $S=S_{a}$ be the Cayley K3 surface. For $X=S^{[2]}$ we use the basis $(h_{1},e,h_{2})$ of $\operatorname{NS}(X)$, in which the BBF form is given by
$q(x,y,z)=4x^{2}+4axz-4z^{2}-2y^{2}.$
For each $k\in\mathbb{Z}$, set
$F_{k}:=a_{2k-1}h_{1}+a_{2k}h_{2}.$
Then $F_{k}^{2}=4$ and, by Saint–Donat’s criterion [26], each divisor $F_{k}$ is very ample. Denote by $\iota_{k}$ the Beauville involution on $X$ induced by the quartic embedding associated to $F_{k}$. In the basis $(h_{1},e,h_{2})$ of $\operatorname{NS}(X)$, the involution $\iota_{0}$ acts by
$\iota_{0}^{*}=\begin{pmatrix}3&2&2a\\ -4&-3&-2a\\ 0&0&-1\end{pmatrix}$.
Define
$P:=\begin{pmatrix}1&0&a\\ 0&1&0\\ a&0&a^{2}+1\end{pmatrix}$.
Since the very ample divisors $F_{k}$ are obtained from $F_{0}$ by the lattice isometry $P^{k}$, the corresponding Beauville involutions are conjugate:
$\iota_{k}^{*}=P^{k}\iota_{0}^{*}P^{-k}.$
Lemma 3.1.
Let $v=(x,y,z)\in\mathbb{Z}^{3}$ satisfy q(v)=2q(v)=-2 with $x>0$. Then $r=z/x\in(\alpha,\beta):=((a-\sqrt{a^{2}+4})/2,(a+\sqrt{a^{2}+4})/2).$
Proof.
Write
$2=q(x,y,z)=4x^{2}\big(1+ar-r^{2}\big)-2y^{2}.$
If a rational number $r$ lies outside the interval $(\alpha,\beta)$, then $x^{2}(1+ar-r^{2})=(x^{2}+axz-z^{2})$ is a negative integer, so $4x^{2}(1+ar-r^{2})\leq-4$, and hence $q(x,y,z)\leq-4-2y^{2}\leq-4$, contradicting the assumption q(x,y,z)=2q(x,y,z)=-2. ∎
By Hassett–Tschinkel’s criterion (Theorem 2.2), the ampleness of a divisor can be tested against classes $\rho$ with $(\rho,\rho)\geq-2$, or against certain exceptional classes with $(\rho,\rho)=-10$ and divisibility two. In our situation, no (10)(-10)-classes occur, and hence the nef cone is obtained by cutting the positive cone by the hyperplanes orthogonal to (2)(-2)-classes.
Theorem 3.2.
The nef cone of $X$ is defined by the hyperplanes orthogonal to the classes in the orbit $\operatorname{Aut}(X)\cdot e$. Moreover,
$\operatorname{Eff}(X)={\rm{Cone}}(\operatorname{Aut}(X)\cdot e).$
Proof.
Suppose that there exists an effective class $v$ with square 2-2 outside the conic hull of the orbit $\operatorname{Aut}(X)\cdot e$. We may assume that $v=(x,y,z)$ spans an extremal ray of the effective cone and that the value $\langle e,v\rangle$ is minimal among all such classes. From $\langle v,e\rangle>0$ and $\langle v,2h_{1}+h_{2}\rangle>0$ it follows that $y<0$ and $x>0$.
The minimality of $\langle e,v\rangle$ implies that
$\langle e,\iota_{k}^{*}v\rangle\geq\langle e,v\rangle$
holds for every integer $k$. Equivalently,
(3.1) $(2a_{2k-1}+a\,a_{2k})x+(a\,a_{2k-1}-2a_{2k})z\geq-2y.$
For the first case $k=0$, the inequality (3.1) reads
$2x+az\geq-2y.$
Since $y<0$, the right-hand side is positive, hence $2x+az>0$, and we may square the inequality to obtain
$(2x+az)^{2}\geq 4y^{2}.$
Substituting q(v)=2q(v)=-2 yields
$2=4x^{2}+4axz-4z^{2}-2y^{2}\geq 4x^{2}+4axz-4z^{2}-\frac{(2x+az)^{2}}{2}.$
Rearranging yields
$4x^{2}+4axz-(a^{2}+8)z^{2}+4\leq 0.$
Dividing by $x^{2}>0$ and setting $r=z/x$, we obtain
$(a^{2}+8)r^{2}-4ar-4\geq\frac{4}{x^{2}}>0.$
Thus $r$ does not lie in the interval
$r\notin I_{0}:=\Biggl[\frac{2a-2\sqrt{2a^{2}+8}}{a^{2}+8},\,\frac{2a+2\sqrt{2a^{2}+8}}{a^{2}+8}\Biggr].$
Since $P(x,y,z)=(x+az,y,ax+(a^{2}+1)z)$, applying $P$ once shifts the coefficients in (3.1) by two indices. In particular, we have
$(2a_{2k-1}+a\,a_{2k})(x+az)+(a\,a_{2k-1}-2a_{2k})(ax+(a^{2}+1)z)$
$\qquad=(2a_{2k-3}+a\,a_{2k-2})x+(a\,a_{2k-3}-2a_{2k-2})z.$
It follows that the transformation $P$ acts on the ratio $r=z/x$ via the Möbius transformation
$f(r)=\frac{a+(a^{2}+1)r}{1+ar}.$
Consequently, for each $k$, the inequality (3.1) imposes the condition $r\notin I_{k}:=f^{k}(I_{0})$.
By Lemma 3.1, the possible values of $r$ lie in the open interval $(\alpha,\beta)$, where $\alpha=\frac{a-\sqrt{a^{2}+4}}{2}$ and $\beta=\frac{a+\sqrt{a^{2}+4}}{2}$. To analyze the action of $f$, define a change of variable
$\psi(r)=\frac{r-\alpha}{\beta-r},$
which maps $(\alpha,\beta)$ bijectively onto $(0,\infty)$. A direct computation shows that
$\psi(f(r))=\beta^{4}\psi(r).$
$r$ must not belong to $I_{k}$ for every $k$. Since $\{\psi(I_{k})\}$ cover $(0,\infty)$, no such $r$ can exist.
Since $e$ is effective, it follows that
${\rm Cone}\bigl(\operatorname{Aut}(X)\cdot e\bigr)\subseteq{\rm Eff}\bigl(X\bigr).$
If the effective cone admits an extremal ray outside the cone generated by this orbit, such a ray would necessarily be rational and generated by a (2)(-2)-class, as no other extremal classes can lie in the closure of the positive cone. However, every (2)(-2)-class that defines a wall of the nef cone is contained in the orbit $\operatorname{Aut}(X)\cdot e$. It follows that no additional extremal rays occur, and hence
${\rm Eff}\bigl(X\bigr)={\rm Cone}\bigl(\operatorname{Aut}(X)\cdot e\bigr).$
Hence, the nef and effective cones are entirely determined by the automorphism group.
Remark $3.$
The proof of Theorem 3.2 extends to the case $a=2$. In this case, the arguments of Lemma 3.1 and Theorem 3.2 fail only in the exceptional case where q(v)=10q(v)=-10 and $|y|=1$. One verifies that
$(v,H+\alpha D)\,(v,H+\beta D)=-8(a^{2}+4)<0.$
Thus the class $v$ takes opposite signs on the two boundary nef rays $H+\alpha D$ and $H+\beta D$, and therefore does not give rise to a wall.
4. Fundamental domain
We work under the same assumptions and notation as in Section 3. The existence of a rational polyhedral fundamental domain in this setting has been shown in general by Amerik and Verbitsky [2]. Our goal in this section is to construct such a fundamental domain explicitly.
We consider three planes
$x-z=0,\qquad z=0,\qquad 3x+2y=0.$
These planes divide the nef-effective cone into four regions:
$\begin{cases}(1)\ x-z<0,\\[2.0pt] (2)\ z<0,\\[2.0pt] (3)\ 3x+2y<0,\\[2.0pt] (4)\ x-z\geq 0,\ z\geq 0,\ 3x+2y\geq 0.\end{cases}$
Indeed, by definition of the nef cone, the inequalities $8x+6y+4z\geq 0$ and $12x+6y-4z\geq 0$ are included. Hence, if $3x+2y<0$, it follows that $0<z<x$.
Including also the inequality $y\leq 0$, the fourth region $(4)$ forms a rational polyhedral cone $\Pi$ bounded by the six hyperplanes
$x-z=0,\quad z=0,\quad 3x+2y=0,\quad 8x+6y+4z=0,\quad 12x+6y-4z=0,\quad y=0,$
and we claim that $\Pi$ is a rational polyhedral fundamental domain of the nef-effective cone of $X$.
Theorem 4.1.
The automorphism group $\operatorname{Aut}(X)\cong\mathbb{Z}_{2}*\mathbb{Z}_{2}*\mathbb{Z}_{2}$ acts discretely on $\operatorname{NS}(X)_{\mathbb{R}}$ preserving the Beauville–Bogomolov–Fujiki form and admits a rational polyhedral fundamental domain $\Pi\subset\operatorname{Nef}^{+}(X)$.
Proof.
Suppose that $p=(x,y,z)$ is a nef $\mathbb{R}$-divisor with $q(p)>0$. We apply the automorphisms iteratively until the point $p$ is contained in the desired region. Let
$A_{1}:=\iota_{1}^{*},\qquad A_{2}:=\iota_{0}^{*},\qquad A_{3}:=\iota_{1}^{*}\iota_{2}^{*}\iota_{1}^{*};$
here $\iota_{k}^{*}$ denotes the Beauville involution attached to $F_{k}$ (see Subsection 2.4, Section 3).
When xz<0x-z<0, applying $A_{1}$ sends $p$ to
$A_{1}p=(5x+2y-2z,\,-6x-3y+2z,\,6x+2y-3z),$
for which (5x+2y2z)(6x+2y3z)>0(5x+2y-2z)-(6x+2y-3z)>0 holds. If, instead, $z<0$, $A_{2}$ yields
$A_{2}p=(3x+2y+2z,\,-4x-3y-2z,\,-z),$
and the third coordinate satisfies z>0-z>0. Finally, if $3x+2y<0$, $A_{3}$ yields
$A_{3}p=(19x+12y,\,-30x-19y,\,10x+6y-z),$
and inequality 3(19x+12y)+2(30x19y)>03(19x+12y)+2(-30x-19y)>0 follows. Thus, depending on the sign conditions of (xz)(x-z), $z$, and $(3x+2y)$, one of the automorphisms $A_{1}$, $A_{2}$, or $A_{3}$ moves $p$ out of its region. We iterate this procedure until $p$ enters the region $(4)$, that is, the rational polyhedral cone $\Pi$ defined by the inequalities above.
At each step of the reduction process, the linear function $f(u,v,w)=2u+v$ decreases strictly:
$f(A_{1}p)-f(p)=2(x-z),\quad f(A_{2}p)-f(p)=2z,\quad f(A_{3}p)-f(p)=2(3x+2y).$
Since each transformation is applied only when the corresponding inequality xz<0x-z<0, $z<0$, or $3x+2y<0$ holds, the value of $f$ strictly decreases at every step. If at any step $f(p)<0$, then $q(p)<0$, contradicting the assumption that $p$ lies in the positive cone.
If the process did not terminate, then at least two of three expressions $f(A_{i}p)-f(p)$ would eventually converge to 0, which would force
$q(x,y,z)$ $=4x^{2}+4xz-4z^{2}-2y^{2}$
$=3\Bigl((x-z)z+(x-z)(3x+2y)+z(3x+2y)\Bigr)$
$\ \ \ \ -\frac{1}{2}\Bigl((x-z)^{2}+z^{2}+(3x+2y)^{2}\Bigr)$
to tend to 0 or below. However, each $A_{i}$ preserves the BBF form, so $q(p)$ remains a fixed positive constant — a contradiction. Therefore, the process must terminate after finitely many steps, and the resulting point lies in region $(4)$, that is, in $\Pi$. Consequently, we have
$\operatorname{Nef}^{+}(X)=\bigcup_{g\in\operatorname{Aut}(X)}g(\Pi).$
We now show that the group $\operatorname{Aut}(X)=\langle A_{1},A_{2},A_{3}\rangle\cong\mathbb{Z}_{2}*\mathbb{Z}_{2}*\mathbb{Z}_{2}$ acts on $\operatorname{Nef}^{+}(X)$ via the translates of $\Pi$ with pairwise disjoint interiors. Let $g=A_{i_{1}}\cdots A_{i_{k}}$ be a reduced word, meaning that no two adjacent generators coincide. Assume for contradiction that $\operatorname{int}(\Pi)\cap\operatorname{int}(g\Pi)\neq\varnothing$. Then there exists a point $p\in\operatorname{int}(\Pi)$ such that $g^{-1}p\in\operatorname{int}(\Pi)$.
By construction, each generator $A_{i_{j}}$ sends every region other than $(i_{j})$ into region $(i_{j})$. Consequently, $A_{i_{k}}^{-1}$ maps region $(4)$ into region $(i_{k})$, then $A_{i_{k-1}}^{-1}$ maps it into region $(i_{k-1})$, and so on. By induction, $g^{-1}\operatorname{int}(\Pi)$ is contained in region $(i_{1})$. However, the interior of region $(4)$ is disjoint from the interiors of all other regions $(i)$. This contradicts the assumption that $g^{-1}p\in\operatorname{int}(\Pi)$.
Therefore, $\operatorname{int}(\Pi)\cap\operatorname{int}(g\Pi)=\varnothing$ whenever $g\neq\mathrm{id}$, and the interiors of distinct translates of $\Pi$ under $\operatorname{Aut}(X)$ are pairwise disjoint. Hence $\Pi$ is a rational polyhedral fundamental domain for the action of $\operatorname{Aut}(X)$ on $\operatorname{Nef}^{+}(X)$, completing the proof. ∎
Figure 4.1. The slice of the nef cone of $S_{1}^{[2]}$ where the coefficient of $h_{1}$ is fixed to be 11
5. Extremal rays of the nef cone for large $n$
In this section, we describe the structure of the nef cone of the Hilbert scheme $X=S_{a}^{[n]}$ and determine its extremal rays when $n$ is sufficiently large.
Let $S=S_{a}$ be the corresponding K3 surface. For each integer $k\in\mathbb{Z}$, recall the divisor classes
$F_{k}=a_{2k-1}h_{1}+a_{2k}h_{2},$
which satisfy
$F_{k}^{2}=4,\qquad F_{k}\!\cdot\!F_{k+1}=2(a^{2}+2).$
For integers $a\geq 1$ and $0\leq t\leq a$, we define the intermediate divisor classes
$E_{k,t}:=\frac{a-t}{a}\,F_{k}+\frac{t}{a}\,F_{k+1}.$
A direct computation shows that
$E_{k,t}^{2}=4\bigl(1+a\,t-t^{2}\bigr),\quad E_{k,t}\!\cdot\!F_{k}=4+2a\,t,\quad E_{k,t}\!\cdot\!F_{k+1}=2(a^{2}+2)-2a\,t.$
Theorem 5.1.
Let $a\geq 1$ and assume
$n\;\geq\;\frac{a^{2}+4a+7}{2}.$
Then the nef cone of $X$ is cut out by the hyperplanes
$(E_{k,t})_{[n]}^{\perp}\quad(k\in\mathbb{Z},\;t=0,\dots,a-1),$
together with $(C_{0})_{[n]}^{\perp}$, where $C_{0}$ is the curve class contracted by the Hilbert–Chow morphism.
Proof.
Let $F=xF_{k}+yF_{k+1}$ be an effective divisor class in $N^{1}(S_{a})$. Consider the matrix
$M_{k}=\begin{pmatrix}a_{2k-1}&a_{2k+1}\\ a_{2k}&a_{2k+2}\end{pmatrix}$.
Since $\det(M_{k})=a$, the sublattice of $N^{1}(S_{a})$ generated by $F_{k}$ and $F_{k+1}$ has index $a$. It follows that $x$ and $y$ must be of the form $x=\frac{s}{a}$ and $y=\frac{t}{a}$, where $s,t\in\mathbb{Z}$. The condition that $F$ has integral coefficients is equivalent to the following pair of congruences
$a_{2k-1}s+a_{2k+1}t\equiv 0,\qquad a_{2k}s+a_{2k+2}t\equiv 0\pmod{a}$.
A particular solution is $(s,t)=(a_{2k+1},-a_{2k-1})$, and every solution is congruent to a multiple of this pair modulo $a$. Hence,
$s+t=a_{2k+1}-a_{2k-1}=a\,a_{2k},$
which is always divisible by $a$, and therefore
$x+y=\frac{s+t}{a}$
is always an integer. It follows that the intermediate classes $E_{k,t}$ are integral.
The vertices of the convex hull of all nontrivial effective divisors are precisely the classes $F_{k}$, and the points $E_{k,t}$ are exactly the lattice points on the edges joining two consecutive vertices $F_{k}$ and $F_{k+1}$.
Define
$C:=an+2a\bigl(t^{2}+t-1\bigr)-2a^{2}t$
and set
$H=(\,C-2a^{2}-4\,)F_{k}\;+\;(\,C+4\,)F_{k+1},\qquad D=-E_{k,t}$
Observe that $H$ is ample if and only if $H^{2}>0$, which is equivalent to
$n>(2t+1)a-2t(t+1)+2+\sqrt{a^{2}+4}$
Assume further that $2n>D^{2}$. Recall that
$\operatorname{CrDiv}(H,D)=\{\,F\in\operatorname{Pic}(S)\mid F\ \text{is effective and}\ H\!\cdot\!F<H\!\cdot\!(-D)\,\}\ \cup\ \{-D\}.$
Let $F=xF_{k}+yF_{k+1}$ be a critical divisor. One computes
$H\!\cdot\!F_{k}-H\!\cdot\!F_{k+1}=4a^{2}(a^{2}+4)>0.$
Thus $H\!\cdot\!F_{k}>H\!\cdot\!F_{k+1}$, and for a fixed sum $m=x+y$, the linear form
$H\!\cdot\!F=xH\!\cdot\!F_{k}+yH\!\cdot\!F_{k+1}$
is minimized when $(x,y)=(0,m)$. In particular, if $x+y\geq 2$, then
$H\!\cdot\!F\geq 2H\!\cdot\!F_{k+1}=H\!\cdot\!E_{k,t}+2a\,(a^{2}+4)\,\Bigl(n+2t^{2}+4t-2-2at-4a\Bigr)\geq H\!\cdot\!E_{k,t}.$
The last inequality holds whenever
$n\geq 2+2(a-t)(t+2).$
Under this condition, the strict inequality $H\!\cdot\!F<H\!\cdot\!(-D)$ cannot occur. Hence, every critical divisor must be of the form $F=E_{k,s}$ for some $s\in\{0,1,\dots,a\}$.
Set
$A:=H\!\cdot F_{k},\qquad B_{0}:=H\!\cdot(F_{k+1}-F_{k}).$
For each $s\in\{0,1,\dots,a\}$, we compare the centers of the divisors $E_{k,s}$ relative to $E_{k,t}$:
$s\!\left(E_{k,s}\right)-s\!\left(E_{k,t}\right)=\frac{(s-t)\bigl(s-t-1\bigr)}{a\,(a^{2}+4)\,\bigl(A-2(s-t)\bigr)}.$
From this expression, we see that $s(E_{k,s})\geq s(E_{k,t})$ for all $s$, and equality holds if and only if $s=t$ or $s=t+1$. Therefore, the only critical divisors with minimal center are
$\operatorname{CrDiv}(H,D)=\{E_{k,t},\,E_{k,t+1}\}.$
The center of the Gieseker wall associated to this pair is given by
$s_{W}=-\frac{1}{2a\,(a^{2}+4)}.$
By Proposition 2.9, the corresponding nef divisor is
$D_{\sigma}$ $\sim-\,s_{W}\,H-D-e$
$=\Biggl[\frac{\,an+2a(t^{2}+t-1)-2a^{2}t-2a^{2}-4\,}{2a(a^{2}+4)}+\frac{a-t}{a}\Biggr]F_{k}$
$\quad+\Biggl[\frac{\,an+2a(t^{2}+t-1)-2a^{2}t+4\,}{2a(a^{2}+4)}+\frac{t}{a}\Biggr]F_{k+1}-e=:D_{k,t}.$
The above construction leads directly to the following description of the nef cone for sufficiently large $n$.
For each $t$, the genus of the curve $E_{k,t}$ is
$g(E_{k,t})=3+2at-2t^{2},$
so the Brill–Noether number $\rho(g,1,n)=2n-g-2$ is nonnegative whenever
$n\geq\frac{5+2at-2t^{2}}{2}.$
In this range, $E_{k,t}$ admits a linear series $g^{1}_{n}$.
For any integral curve $C\subset S$, the intersection with a divisor $xh_{1}+ye+zh_{2}$ is given by
$(xh_{1}+y\,e+zh_{2})\cdot C_{[n]}=(xh_{1}+zh_{2})\!\cdot\!C+y(n+g(C)-1).$
Applying this to $D_{k,t}$ and $C=E_{k,u}$ yields
$D_{k,t}\cdot(E_{k,u})_{[n]}=2(u-t)(u-t-1).$
Thus
$D_{k,t}\cdot(E_{k,t})_{[n]}=D_{k,t}\cdot(E_{k,t+1})_{[n]}=0,$
and $D_{k,t}$ is positive on all other $(E_{k,u})_{[n]}$. It follows that $D_{k,t}$ contracts precisely the two curves $(E_{k,t})_{[n]}$ and $(E_{k,t+1})_{[n]}$, and therefore spans an extremal nef ray of $X=S_{a}^{[n]}$.
The supporting hyperplanes associated to these curves are
$(E_{k,t})_{[n]}^{\perp}=\Bigl\{\,xh_{1}+y\,e+zh_{2}\ \Big|\ (xh_{1}+zh_{2})\!\cdot\!E_{k,t}+y\Bigl(n+\frac{E_{k,t}^{2}}{2}\Bigr)=0\Bigr\},$
and analogously for $(E_{k,t+1})_{[n]}^{\perp}$. Since $D_{k,t}$ lies in the intersection of these two hyperplanes and they are both supporting, the ray $\mathbb{R}_{\geq 0}[D_{k,t}]$ is an extremal ray of $\operatorname{Nef}(X)$.
Next, we consider the limiting behavior of the divisors $F_{k}=a_{2k-1}h_{1}+a_{2k}h_{2}$. As $k\to+\infty$ and $k\to$-$\infty$, the ratios $r_{k}=a_{2k}/a_{2k-1}$ converge to
$\beta=\frac{a+\sqrt{a^{2}+4}}{2},\qquad\alpha=\frac{a-\sqrt{a^{2}+4}}{2}.$
Hence, the normalized equations of the supporting hyperplanes $(F_{k})_{[n]}^{\perp}$ converge to
$(4+2\beta)x+(2-4\beta)z+(n+2)y=0,\qquad(4+2\alpha)x+(2-4\alpha)z+(n+2)y=0.$
On the other hand, the curve class $C_{0}$ satisfies $e\cdot C_{0}=-1$ and $h_{1}\cdot C_{0}=h_{2}\cdot C_{0}=0$, so
$C_{0}^{\perp}=\{\,xh_{1}+ye+zh_{2}\mid y=0\,\}.$
Intersecting $C_{0}^{\perp}$ with the two limiting hyperplanes gives the remaining two extremal rays of $\operatorname{Nef}(X)$, generated by $h_{1}+\beta h_{2}$ and $h_{1}+\alpha h_{2}$. These rays are irrational and lie in the closure of $\Lambda$.
Finally, the argument requires the following numerical conditions:
$n$ $>(2t+1)a-2t(t+1)+2+\sqrt{a^{2}+4},$
$n$ $>E_{k,t}^{2}/2,$
$n$ $\geq 2+2(a-t)(t+2),$
$n$ $\geq\frac{5+2at-2t^{2}}{2}.$
The bound
$n\geq\frac{a^{2}+4a+7}{2}$
arises from the third inequality
$n\geq 2+2(a-t)(t+2)$
when evaluated at $t=\lfloor(a-1)/2\rfloor$. ∎
In the proof of the theorem we obtain another description:
Corollary 5.2.
The nef cone of $X$ coincides with the closure of the cone
$\Lambda\;:=\;\mathrm{Cone}\big\langle D_{k,t}\mid k\in\mathbb{Z},\;t=0,\dots,a-1\big\rangle.$
Note that the Mori cone is dual to the nef cone. It is generated by the curve classes $(E_{k,t})_{[n]}$, the exceptional curve $C_{0}$, and the two limiting classes $C_{+}$ and $C_{-}$ obtained as limits of $(F_{k})_{[n]}$ as $k\to\pm\infty$. These limiting classes are dual to the irrational supporting hyperplanes described above, and hence the corresponding extremal rays of the Mori cone are also irrational.
Example 5.3.
In the case $a=1$ and $n\geq 6$, the nef cone of $S_{1}^{[n]}$ is determined in a simple way by the hyperplanes $(F_{k})_{[n]}^{\perp}$ and $C_{0}^{\perp}$, where $F_{k}=a_{2k-1}h_{1}+a_{2k}h_{2}$, with coefficients given by the classical Fibonacci sequence. Moreover, the extremal rays are $D_{k,0}=\dfrac{n+2}{10}\,(F_{k}+F_{k+1})-e$, $h_{1}+\alpha h_{2}$, and $h_{1}+\beta h_{2}$.
Remark 5.4.
For smaller values of $n$, the wall structure of the nef cone can become more complicated and is not always determined solely by the divisors $E_{k,t}$. In such cases, new types of walls may appear, and more refined techniques are required to describe the nef and Mori cones.
Figure 5.1. The slice of the nef cone of $S_{1}^{[6]}$ where the coefficient of $h_{1}$ is fixed to be 11